OpenStudy (seniaa.tapia):

find the first,fourth snd tenth terms of the Arithmetic sequence A(n)=-11+(n-1)(4)

myininaya (myininaya):

to find the first term replace n with 1 to find the fourth term replace n with 4 and so on...

myininaya (myininaya):

For example, if I wanted the sixth term I replace n with 6 like so:\[A(6)=-11+(6-1)(4) \\ A(6)=-11+5(4) \\ A(6)=-11+20 \\ A(6)=9\] the 6th term of the sequence A is 9.